# Inertial Magnetic SLAM Systems Using Low-Cost Sensors

### v1

@@ 3.2 The Magnetic Field-Aided Inertial Navigation System
Here $f^{\text{\tiny{ins}}}(\cdot)$ denotes the inertial subsystem dynamics, $f^{\theta}(\cdot)$ denotes the magnetic subsystem function. The exact form of these functions is presented in [11]. Moreover, $\tilde{u}_{k}=[\tilde{a}_{k}^{\top}\;\tilde{\omega}_{k}^{\top}]^{\top}$ denotes the IMU measurements, where $\tilde{a}_{k}\in\mathbb{R}^{3}$ and $\tilde{\omega}_{k}\in\mathbb{R}^{3}$ denote the specific force and angular velocity, respectively. Furthermore, $w_{k}=[(w_{k}^{\text{\tiny{ins}}})^{\top}\;(w_{k}^{\theta})^{\top}]^{\top}$ denotes the process noise, where $w_{k}^{\text{\tiny{ins}}}\in\mathbb{R}^{12}$ and $w_{k}^{\theta}\in\mathbb{R}^{8}$ denote the process noise of the inertial and magnetic subsystems, respectively. The measurement vector $y_{k}\in\mathbb{R}^{3N}$ denotes the magnetometer measurements from an array of $N$ magnetometers, where $r^{(i)}$ denotes the position of the $i$-th magnetometer in the array expressed in the body frame. $\Phi(\cdot)$ denotes the regressor matrix in (6a). Finally, $e_{k}\in\mathbb{R}^{3N}$ denotes the white Gaussian measurement noise.
Figure 4: The sensor board used in the magnetic field-aided inertial navigation system. It has 30 PNI RM3100 magnetometers and an Osmium MIMU 4844 IMU mounted on the bottom side. (a) Front side of the sensor board showing the magnetometer array. (b) Back side of the sensor board showing the IMU.
@@ 3.3 Summary
Both the magnetic field SLAM system and the MAINShave their own advantages and limitations. The magnetic field SLAM system has a bounded positioning error in mapped regions. However, its reliance on the low-drift visual odometry limits its usability. The MAINS, on the other hand, can provide low-drift inertial navigation, but it only maintains a local magnetic field map and, thus, cannot provide positioning with bounded error. In the next section, we propose a loosely-coupled magnetic field SLAM system that combines the advantages of both systems while mitigating their limitations.
4 The Loosely-Coupled IM-SLAMSystem
A loosely coupled IM-SLAM system can be regarded as a cascaded architecture composed of a modified MAINSand a modified magnetic field SLAM system, as illustrated in Fig. 5. The dependence on low-drift visual odometry data is eliminated by replacing it with odometry provided by the modified MAINSsystem. Furthermore, the measurement model in the magnetic field SLAM system is adapted to support magnetometer arrays, rather than a single magnetometer, as well as an additional barometer. The barometer supplies altitude information, enabling the SLAM system to achieve improved 3D trajectory estimation.
Figure 5: An overview of the loosely-coupled IM-SLAMsystem. The system is a cascade of a modified MAINS, an odometry data conversion module, and a modified magnetic field SLAM system.
4.1 Computing Odometry Data with the MAINS
The original MAINS [11] estimates the pose at the current timestep; therefore, it cannot be used directly to compute odometry data. One way to make the MAINS support odometry data computation is to augment the state vector of MAINSto include the complete state vector (or the pose) at the past and compute the difference between the pose at the pase and the current. A detailed explanation of this modification is presented next. The augmented state vector is defined as
$x_{k}^{a}$ $\triangleq[x_{k}^{\text{\tiny{mains}}\top}\;x_{k}^{\text{\tiny{past}}\top}]^{\top}$ (15a)
$x_{1}^{\text{\tiny{past}}}$ $\triangleq x_{1}^{\text{\tiny{mains}}}$ (15b)
$x_{k}^{\text{\tiny{past}}}$ $\triangleq f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}}),\quad k\geq 2$ (15c)
$f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}})$ $=\begin{cases}x_{k-1}^{\text{\tiny{past}}}&nK+3\leq k\leq(n+1)K+1\\ x_{k-1}^{\text{\tiny{mains}}}&k=nK+2\end{cases}$ (15d)
$n$ $=0,1,2,\cdots$ (15e)
Here $K\in\mathbb{N}$ denotes the number of time steps within the time interval for which the odometry data is computed. The state-space model of the modified MAINSis then
$x_{k+1}^{a}=f^{a}(x_{k}^{a},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{mains}}}(x_{k}^{\text{\tiny{mains}}},\tilde{u}_{k},w_{k})\\ f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}})\end{bmatrix}$ (16a)
@@ 4.1 Computing Odometry Data with the MAINS
$H^{a}$ $=\begin{bmatrix}0_{3\times 16}&\Phi(r^{(1)})&0_{3\times 16}\\ \vdots&\vdots&\vdots\\ 0_{3\times 16}&\Phi(r^{(N)})&0_{3\times 16}\end{bmatrix}$. (16c)
The modified MAINS can compute odometric data between the time step $i=nK+1$ and $j=(n+1)K+1$ using the posterior state estimate at time step $j$ as
$\tilde{u}^{o}_{i,j}$ $=\begin{bmatrix}\Delta p_{i,j}\\ \Delta q_{i,j}\end{bmatrix}=\begin{bmatrix}\hat{p}_{j}-\hat{p}_{i}\\ \hat{q}_{i}^{\ast}\otimes\hat{q}_{j}\end{bmatrix}$ (17a)
where $(\hat{\cdot}_{i})$ and $(\hat{\cdot}_{j})$ denote the smoothed and filtered quantity at time $i$ and $j$, respectively, and $(\cdot)^{\ast}$ denotes conjugated quaternion. The covariance of the odometry data can be approximated by
$\text{Cov}(\tilde{u}_{i,j}^{o})\approx A_{i,j}\;\text{Cov}(\hat{p}_{i},\hat{p}_{j},\hat{q}_{i},\hat{q}_{j})\;A_{i,j}^{\top}$ (17b)
@@
where $\text{Cov}(\cdot)$ denotes the covariance. Note the covariance of the quaternion is defined in the tangent space of $S^{3}$, see [27, p. 54] for details.
@@ 4.2 The Modified Magnetic Field SLAM System
The modified magnetic field SLAM system uses the odometric data provided by the modified MAINS, magnetic field measurements from the sensor array, and barometer measurements. The state vector of this system and the dynamics remain the same as in Section 3.1. The measurement equation is modified to
$y_{k}=h^{\text{\tiny{array-slam}}}(x_{k}^{\text{\tiny{SLAM}}})+e_{k}$ (18a)
@@
$h^{\text{\tiny{array-slam}}}(x_{k}^{\text{\tiny{SLAM}}})=\begin{bmatrix}R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(1)}\big)\\ R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(2)}\big)\\ \vdots\\ R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(N)}\big)\end{bmatrix}\eta_{k}.$ (18b)
Here $h^{\text{\tiny{array-slam}}}(\cdot)$ denotes the modified measurement function, which takes into account the multiple sensors and the geometric configuration of the sensor array. This measurement equation is augmented when barometric measurements are available, i.e.,
$y_{k}^{b}=p_{z,k}+e_{k}^{b}$ (19)
where $p_{z,k}\in\mathbb{R}$ denotes the altitude (the third component of the position vector), and $e_{k}^{b}\in\mathbb{R}$ denotes the barometer measurement noise, modeled as Gaussian white noise.
@@ 4.3 Summary
The loosely-coupled IM-SLAM system combines the advantages of both the magnetic field SLAM system and the MAINS. It is a self-contained system that can provide bounded-error positioning in mapped areas without the need for external low-drift odometry. However, there are two main limitations of this system. First, the magnetometer measurements are used twice: first in the modified MAINSand then in the modified magnetic field SLAM system. Second, the odometric data provided by the modified MAINSis correlated over time, which violates the assumption of uncorrelated odometric data in the magnetic field SLAM system. These two limitations can potentially degrade the performance of the loosely-coupled IM-SLAM system. To address these limitations, we propose a tightly-coupled IM-SLAM system in the next section.
5 The Tightly-Coupled IM-SLAMSystem
The tightly-coupled IM-SLAM system integrates the inertial, barometric, and magnetic sensors into a unified SLAM framework (see Fig. 6). In this framework, the state vector encompasses the full set of states, including the inertial states, local and global magnetic field model coefficients. The system dynamics and measurement equation are jointly formulated to leverage all available information to achieve better performance. In the following, the state-space model of this system is presented.
Figure 6: An overview of the tightly-coupled IM-SLAMsystem. This system consists of an inertial navigation system, a barometer model, and a local and global magnetic field model. These components are integrated into a unified SLAMframework.
@@ 5.1 State Dynamics
The state vector of the tightly-coupled SLAM system at time step $k$ is defined as
$x_{k}\triangleq[p_{k}^{\top}\;v_{k}^{\top}\;q_{k}^{\top}\;b_{a,k}^{\top}\;b_{g,k}^{\top}\;\theta_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$ (20)
The state transition model is
$x_{k+1}=f(x_{k},\tilde{u}_{k},w_{k})$ (21)
@@
$f(x_{k},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{ins}}}(x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\text{\tiny{ins}}}_{k})\\ f^{\theta}(\theta_{k},x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\theta}_{k})\\ \eta_{k}\end{bmatrix}$. (22)
@@ 5.2 Measurement Equations
Since the state vector includes both the local and global magnetic field model coefficients, the magnetometer array’s measurements can be expressed in terms of both models. To achieve IM-SLAM, one can use the measurement equations in an alternating manner from either the MAINSor the magnetic field SLAM system. A natural choice is to use the computationally efficient measurement equation from the MAINSmost of the time, while occasionally using the more complex measurement equation from the magnetic field SLAM system. In this way, the system can benefit from the low-drift inertial navigation provided by the MAINSwhile still leveraging the global consistency of the magnetic field SLAM system. In the proposed tightly-coupled IM-SLAM, the measurement equation of the global model, which is based on (18), is further simplified to reduce computational complexity. The basic idea is that the magnetic field at any sensor’s location can be viewed as a superimposition of the magnetic field in the center of the array and some local magnetic field variation. This results in
@@
To sum up, there are two measurement equations for magnetometer array measurements: the equation used by the MAINSand the equation (23). They are used in an alternating manner: the former is employed at every update step except once $D$ samples, when the second model is used instead. That is,
@@ 6 Experimental Results
To evaluate the performance of the proposed loosely-coupled and tightly-coupled IM-SLAMsystems, we conducted experiments in an indoor environment using a custom-built sensor array, see Fig. 4. The sensor array is equipped with an IMU, a barometer, and an array of magnetometers.
In each experiment, a person carried the sensor array and started moving in a room equipped with a motion capture system for several tens of seconds. The person then walked out of the room and explored the surrounding area for several minutes before returning to the room. In total, three types of trajectories were recorded: a trajectory with long straight segments, a trajectory with square loops, and a trajectory with spiral upward segments (climbing a spiral staircase). The characteristics of the datasets are summarized in TABLE 1.
The IMUand magnetometers were sampled at 100 Hz, while the barometer was sampled at 10 Hz. When available, the motion capture system provided ground truth poses at 100 Hz. In the loosely-coupled IM-SLAM system, the odometry data was internally computed at $5$ Hz (K=20K\!=\!20). In the tightly-coupled SLAM system, the magnetic field model switching time was set to 1 second (D=100D\!=\!100), the length scale of the approximated GPmodel $l_{\text{SE}}$ was set to 0.7 meters. More details about the parameter settings can be found in the code repository.
The standard error-state extended Kalman filter (EKF)is used in all systems for state estimation; interested reader are referred to [27] for details.
@@
6.1 Comparison with the MAINS
In the first experiment, the loosely-coupled and tightly-coupled IM-SLAM systems were compared with MAINS. Note that MAINS was modified in the same way as in Section 3.1 to integrate barometer measurements.
All three systems used the pose from the motion capture system as additional measurements when the sensor array was in the room at the beginning of the trajectory. This is to stablize the IMU bias estimates, which is important for good positioning performance. During the later parts of the trajectory, the systems relied solely on measurements from the onboard sensors for positioning.
Three examples of the trajectories estimated by the three systems are shown in Fig. 7, Fig. 8, and Fig. 9, respectively. The quantitative results are summarized in TABLE 2 and TABLE 3, where the horizontal error, vertical error, and yaw error at the end of the trajectories are reported. The horizontal error is defined as the Euclidean distance between the estimated position and the ground truth position projected onto the horizontal plane, while the vertical error is defined as the absolute difference between the estimated altitude and the ground truth altitude. The yaw error is defined as the absolute difference between the estimated yaw angle and the ground truth yaw angle.
It can be seen from the figures that all systems can track the general shape of the trajectories. The noticeable difference between the tightly coupled IM-SLAMand other systems is that the estimated trajectories by the former have more jerky segments, which occur when the systems switch between the two measurement models; see (25). This phenomenon could be attributed to two different modeling of the magnetic field models: the model in the first row of (25) is parameterized with the position of the body frame in the navigation frame while the model in the second row is parameterized with the (local) positions of the magnetometers in the body frame. This implies that the former allows direct observation of the position while the latter does not [28]. The sudden jump observed on the estimated trajectory is a sign of gaining positional information when switching from the latter to the former model. Another observation is that the trajectories estimated by the SLAM systems during the long straight segments are more concentrated than those estimated by MAINS (see Fig. 7), which indicates that the SLAM systems can better mitigate the drift during these segments. Moreover, from TABLE 2 and TABLE 3, it can be seen that the tightly-coupled IM-SLAM system outperforms both MAINS and the loosely-coupled IM-SLAM system in most cases.
Table 2: Horizontal (Vertical) Error at the end of the Trajectories with a Barometer. Unit: meter

### v2

@@ 3.2 The Magnetic-Field-Aided Inertial Navigation System
Here $f^{\text{\tiny{ins}}}(\cdot)$ denotes the inertial subsystem dynamics, $f^{\theta}(\cdot)$ denotes the magnetic subsystem function. The exact form of these functions are presented in [11]. Moreover, $\tilde{u}_{k}=[\tilde{a}_{k}^{\top}\;\tilde{\omega}_{k}^{\top}]^{\top}$ denotes the IMU measurements, where $\tilde{a}_{k}\in\mathbb{R}^{3}$ and $\tilde{\omega}_{k}\in\mathbb{R}^{3}$ denote the specific force and angular velocity, respectively. Furthermore, $w_{k}=[(w_{k}^{\text{\tiny{ins}}})^{\top}\;(w_{k}^{\theta})^{\top}]^{\top}$ denotes the process noise, where $w_{k}^{\text{\tiny{ins}}}\in\mathbb{R}^{12}$ and $w_{k}^{\theta}\in\mathbb{R}^{8}$ denote the process noise of the inertial and magnetic subsystems, respectively. The measurement vector $y_{k}\in\mathbb{R}^{3N}$ denotes the magnetometer measurements from an array of $N$ magnetometers, where $r^{(i)}$ denotes the position of the $i$-th magnetometer in the array expressed in the body frame. $\Phi(\cdot)$ denotes the regressor matrix in (6a). Finally, $e_{k}\in\mathbb{R}^{3N}$ denotes the white Gaussian measurement noise.
Figure 1: The sensor board used in MAINS. It has 30 PNI RM3100 magnetometers and an Osmium MIMU 4844 IMU mounted on the bottom side. (a) Front side of the sensor board showing the magnetometer array. (b) Back side of the sensor board showing the IMU.
@@ 3.3 Summary
Both the magnetic field SLAM system and the MAINS have their own advantages and limitations. The magnetic field SLAM system has a bounded positioning error when in previously mapped areas. However, its reliance on the low-drift odometry limits its usability. MAINS, on the other hand, can provide low-drift inertial navigation, but it only maintains a local magnetic field map; thus, positioning errors grow without bound. In the next section, we propose a loosely coupled magnetic field SLAM system that combines the advantages of both systems while mitigating their limitations.
4 The Loosely Coupled IM-SLAM System
The loosely coupled IM-SLAM system integrates a modified MAINS [11] with a modified magnetic field SLAM system [7] in a cascaded configuration. It operates across two distinct time scales, following a common paradigm in visual SLAM systems, i.e., a high-rate odometry process and a low-rate optimization process [31].
In our implementation, the modified MAINS runs at the IMU rate and is indexed by $k$, whereas the modified magnetic field SLAM system operates at a rate downsampled by a factor of $K$ and is indexed by $n$. The relationship between the two time indices is given by $k=1+nK$.
This design reduces the overall computational load, as the magnetic field SLAM subsystem is computationally intensive, typically having a state dimension an order of magnitude larger than that of the modified MAINS.
4.1 Computing Odometry Data with MAINS
The original MAINS [11] estimates only the current pose; therefore, it cannot be used directly to compute odometry data. One approach to enable odometry data computation in MAINS is to augment its state vector with a past state (or pose), and compute the relative motion between the past and current poses. A detailed explanation of this modification is presented next.
The augmented state vector is defined as
$\mathbf{x}^{a}_{k}\triangleq\begin{bmatrix}\mathbf{x}^{\text{\tiny{mains}}}_{k}\\ \mathbf{x}^{\text{\tiny{past}}}_{k}\end{bmatrix},\qquad\mathbf{x}^{\text{\tiny{past}}}_{1}\triangleq\mathbf{x}^{\text{\tiny{mains}}}_{1}.$ (14a)
Here $x_{k}^{\text{\tiny{past}}}$ denotes the past state at the start of the interval over which the odometry data is computed. The state dynamics is defined as
$x^{a}_{k+1}$ $=f^{a}\!\left(x^{a}_{k},\tilde{u}_{k},w_{k}\right)$ (14b)
$f^{a}\!\left(x^{a}_{k},\tilde{u}_{k},w_{k}\right)$ $\triangleq\begin{bmatrix}f^{\text{\tiny{mains}}}\!\left(x^{\text{\tiny{mains}}}_{k},\tilde{u}_{k},w_{k}\right)\\ \begin{cases}x^{\text{\tiny{mains}}}_{k},&\text{if }k\in\mathcal{K}_{\mathrm{upd}}\\ x^{\text{\tiny{past}}}_{k},&\text{otherwise}\end{cases}\end{bmatrix}$. (14c)
Here, $\mathcal{K}_{\mathrm{upd}}=\{1,\,1+K,\,1+2K,\dots\}$ denotes the set of time steps at which odometry data are computed. The state-space model of the modified MAINS is then
$x_{k+1}^{a}$ $=f^{a}(x_{k}^{a},\tilde{u}_{k},w_{k})$ (15a)
@@ 4.1 Computing Odometry Data with MAINS
$H^{a}$ $=\begin{bmatrix}H^{\text{\tiny{mains}}}\quad 0\end{bmatrix}$. (15c)
The error-state Kalman filter [27] is applied here for state estimation. To avoid the double use of the magnetometer measurements in the modified magnetic field SLAM system presented in the next subsection, no measurement update is performed at any time in $\mathcal{K}_{\mathrm{upd}}$.
The modified MAINS computes the odometric data between the time step $i=1+nK$ and $j=1+(n+1)K$ using the state estimate at time step $j$ as
$\tilde{u}^{o}_{n}$ $=\begin{bmatrix}\Delta p_{i,j}\\ \Delta q_{i,j}\end{bmatrix}=\begin{bmatrix}\hat{p}_{j}-\hat{p}_{i}\\ \hat{q}_{i}^{\ast}\otimes\hat{q}_{j}\end{bmatrix}$ (16a)
where $(\hat{\cdot}_{i})$ and $(\hat{\cdot}_{j})$ denote the smoothed and filtered quantities at time $i$ and $j$, respectively, and $(\cdot)^{\ast}$ denotes the quaternion conjugation. The covariance of the odometry data can be approximated by
$\text{Cov}(\tilde{u}_{n}^{o})\approx A_{i,j}\;\text{Cov}(\hat{p}_{i},\hat{p}_{j},\hat{q}_{i},\hat{q}_{j})\;A_{i,j}^{\top}$ (16b)
@@
where $\text{Cov}(\cdot)$ denotes the covariance. Note that the covariance matrix on the right hand side of (16b) can be extracted from the covariance matrix computed by the filter, and the covariance of the quaternion is defined in the tangent space of $S^{3}$, see [27, p. 54] for details.
@@ 4.2 The Modified Magnetic Field SLAM System
The modified magnetic field SLAM system uses the odometric data $\tilde{u}^{o}_{n}$ provided by the modified MAINS, magnetic field measurements from the sensor array at time steps in $\mathcal{K}_{\mathrm{upd}}$, and barometer measurements. The state vector of this system and the dynamics remain the same as in Section 3.1. The measurement equation is modified to
$y_{n}=h^{\text{\tiny{array-slam}}}(x_{n}^{\text{\tiny{SLAM}}})+e_{n}$ (17a)
@@
$h^{\text{\tiny{array-slam}}}(x_{n}^{\text{\tiny{SLAM}}})=\begin{bmatrix}R_{n}^{\top}\nabla\Psi\big(p_{n}+R_{n}r^{(1)}\big)\\ R_{n}^{\top}\nabla\Psi\big(p_{n}+R_{n}r^{(2)}\big)\\ \vdots\\ R_{n}^{\top}\nabla\Psi\big(p_{n}+R_{n}r^{(N)}\big)\end{bmatrix}\eta_{n}.$ (17b)
Here $h^{\text{\tiny{array-slam}}}(\cdot)$ denotes the modified measurement function, which takes into account the multiple sensors and the geometric configuration of the sensor array.
This measurement equation is augmented when barometric measurements are available at time steps in $\mathcal{K}_{\mathrm{upd}}$, i.e., $y_{n}^{b}=p_{z,n}+e_{n}^{b}$, where $p_{z,n}\in\mathbb{R}$ denotes the altitude (the third component of the position vector), and $e_{n}^{b}\in\mathbb{R}$ denotes the barometer measurement noise, modeled as Gaussian white noise.
@@ 4.3 Summary
The loosely coupled IM-SLAM system is a straightforward implementation based on the magnetic field SLAM system and MAINS. However, the main issue is that the odometric data provided by the modified MAINS is correlated over time, which violates the assumption of uncorrelated odometric data in the magnetic field SLAM system. Notably, the temporal correlations in the odometry data are not unique to the proposed system. Visual and visual–inertial odometry estimates are subject to the same effect, since successive motion estimates are computed from overlapping measurements [29]. Nevertheless, this limitation can potentially degrade the performance of the loosely coupled IM-SLAM system. To address this limitation, we propose a tightly coupled IM-SLAM system in the next section.
5 The Tightly Coupled IM-SLAM System
The tightly coupled IM-SLAM system integrates the inertial, barometric, and magnetic sensors into a unified SLAM framework. In this framework, the state vector encompasses the full set of states, including the inertial navigation states, and the local and global magnetic field model coefficients. The system dynamics and measurement equation are jointly formulated to leverage all available information to achieve better performance. Next, the state-space model of this system is presented.
@@ 5.1 State Dynamics
The state vector of the tightly coupled SLAM system at time step $k$ is defined as $x_{k}\triangleq[p_{k}^{\top}\;v_{k}^{\top}\;q_{k}^{\top}\;b_{a,k}^{\top}\;b_{g,k}^{\top}\;\theta_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$, and the state transition model is
$x_{k+1}=f(x_{k},\tilde{u}_{k},w_{k})$ (18a)
@@
$f(x_{k},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{ins}}}(x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\text{\tiny{ins}}}_{k})\\ f^{\theta}(\theta_{k},x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\theta}_{k})\\ \eta_{k}\end{bmatrix}$. (18b)
@@ 5.2 Measurement Equations
Since the state vector includes both the local and global magnetic field model coefficients, the magnetometer array’s measurements can be expressed in terms of both models. To exploit their complementary advantages, one can use the measurement equations in the MAINS or the magnetic field SLAM system in an alternating manner. A natural choice is to use the computationally efficient measurement equation from the MAINS most of the time, while occasionally using the more complex measurement equation from the magnetic field SLAM system. In this way, the system can benefit from the low-drift inertial navigation provided by the MAINS while still leveraging the global consistency of the magnetic field SLAM system. In the proposed tightly coupled IM-SLAM, the measurement equation of the global model, which is based on (17), is further simplified to reduce computational complexity. The basic idea is that the magnetic field at any sensor’s location can be viewed as a superposition of the magnetic field in the center of the array and some local magnetic field variation. This results in
@@
To sum up, there are two measurement equations for magnetometer array measurements: the equation (15c) used by MAINS and the equation (19). They are used in an alternating manner: the former is employed at every update step except once every $D$ samples, when the second model is used instead. That is,
@@
5.3 Summary
The tightly coupled IM-SLAM system augments the INS state with both local and global magnetic field model coefficients, performing measurement updates with these models in an alternating fashion. By estimating the navigation state in a single step, the system bypasses the odometry data computation described in 4.1, thereby resolving the temporal correlation issues inherent in loosely coupled IM-SLAM. However, the computational demand is significantly higher; the state dimension is comparable to that of the modified magnetic field SLAM system and operates at the IMU rate, even though the complex measurement update in (19) is invoked less frequently.
@@ 6 Experimental Results
To evaluate the performance of the proposed loosely coupled and tightly coupled IM-SLAM systems, we conducted experiments in an indoor environment using a custom-built sensor array, see Fig. 1. The sensor array is equipped with an IMU, a barometer, and 30 magnetometers.
In each experiment, a person carried the sensor array and started moving in a room equipped with a motion capture system for several tens of seconds. The person then walked out of the room and explored the surrounding area for several minutes before returning to the room. In total, three types of trajectories were recorded: a trajectory with long straight segments, a trajectory with square loops, and a trajectory with spiral upward segments (climbing a spiral staircase). They are named Long Corridor (LC), Corridor (C), and Spiral Staircase (SS). The characteristics of these datasets are summarized in TABLE 1.
The IMU and magnetometers were sampled at 100 Hz, while the barometer was sampled at 10 Hz. When available, the motion capture system provided ground truth poses at 100 Hz. In the loosely coupled IM-SLAM system, the odometry data was internally computed every second (K=100K\!=\!100). In the tightly coupled SLAM system, the magnetic field model switching time was set to 1 second (D=100D\!=\!100), the length scale of the approximated GP model $l_{\text{SE}}$ was set to 0.7 to 1 meter depending on the datasets. The choice of $D$ and $K$ was based on experimental results. Too large a value would cause SLAM systems to degrade to MAINS, while too small a value would cause the filter to diverge. A study on the effect of model switching time is conducted in Section 6.4.
@@
6.1 Comparison with MAINS
In the first experiment, the loosely coupled and tightly coupled IM-SLAM systems were compared with MAINS. Note that MAINS was modified in the same way as in Section 3.1 to integrate barometer measurements.
All three systems used the pose from the motion capture system as additional measurements at the beginning of the trajectory. This is to stabilize the IMU bias estimates, which is important for good positioning performance [30, Ch. 22]. For many applications, such as first-responder localization, one can expect to have similar aiding at the start, such as GPS positioning. During the later parts of the trajectory, the systems relied solely on measurements from the onboard sensors for positioning.
Three examples of the trajectories estimated by the three systems are shown in Fig. 2, Fig. 3, and Fig. 4, respectively. The quantitative results are summarized in TABLE 2, where the horizontal error and vertical error at the end of the trajectories are reported. The horizontal error is defined as the Euclidean distance between the estimated position and the ground truth position projected onto the horizontal plane, while the vertical error is defined as the absolute difference between the estimated altitude and the ground truth altitude.
It can be seen from the figures that all systems can track the general shape of the trajectories. The noticeable difference between the IM-SLAM systems and MAINS is that the estimated trajectories by the former have more jerky segments, which occur when the loosely coupled system uses (17) and the tightly coupled system uses (19) to update position estimates. These jerky segments are signs of gaining absolute positional information because the measurement equations (17) and (19) are parameterized with the absolute position of the body frame, while (15c) does not allow observation of absolute positions [28]. One way to smooth the trajectory is to perform an extended Kalman filter (EKF) smoothing procedure following the measurement updates concerning (17) and (19). For the tightly coupled system, one can also introduce a pseudo measurement and perform measurement update together with (19), see Section 6.5.
Another observation is that the loosely coupled system and the tightly coupled system performed noticeably differently on Dataset Long corridor-2, where a trajectory with two long segments (branches) and a trajectory with two overlapping segments are reported by them respectively. Given that similar magnetic field magnitudes were observed on these two branches of the trajectory (the loosely coupled system), it is very likely that they were, in fact, overlapping, indicating that the tightly coupled system performs better in this case. Moreover, from TABLE 2, it can be seen that the tightly coupled IM-SLAM system outperforms both MAINS and the loosely coupled IM-SLAM system in most cases.
Table 2: Horizontal and Vertical Error at the End of the Trajectories with a Barometer. Unit: meter